OpenStudy (safyousuff97):

Task 3—Graphing Rational Functions Create a rational function with a linear binomial in both the numerator and denominator. Part 1. Graph your function using technology. Include the horizontal and vertical asymptotes and the x- and y-intercepts on your graph. Label the asymptotes and intercepts. Part 2. Show all work to identify the vertical asymptote, the x-intercepts, and the y-intercept.

OpenStudy (safyousuff97):

I used this function y=x−2x−3 vertical asymptote x = 3 horizontal asymptote y = 0 x intercept 2, y intercept 2/3

OpenStudy (safyousuff97):

can somebody draw the graph pls

OpenStudy (safyousuff97):

@FaireGaming

OpenStudy (varaug):

ill try gimmie a sec.

OpenStudy (safyousuff97):

ok thank you

OpenStudy (fairegaming):

Im looking and this and I am like Whaaaaaa. XD can't you turn this into standard form, sorry I am dumb af

OpenStudy (safyousuff97):

its ok

OpenStudy (mathmale):

Your "y=x−2x−3 " is a polynomial, NOT a rational function. Please review "rational function" and create such a function.

OpenStudy (varaug):